Solve by completing the square.
x^2 + 3/2x = 3
I am stumped because the problem isn't set up the way I've been accustomed to. Can you help me?

Answer by edjones(8007)

x^2+ 3x/2=3
x^2+ 3x/2+ 9/16=3+ 9/16 (multiply 3/2 * 1/2 and square the result to complete the square.
(x+ 3/4)^2=57/16
x+ 3/4=+-sqrt(57)/4 Sqrt of each side.
x=-3/4+-sqrt(57)/4
x=-(3-sqrt(57))/4, x=-(3+sqrt(57))/4
.
Ed